OpenStudy (anonymous):

∫ e^(2x)*(1+e^(2x))^3 dx u=?

OpenStudy (anonymous):

Wolfram Alpha buddy

OpenStudy (anonymous):

take u = 1 + e^(2x)

OpenStudy (anonymous):

let \(\large u=1+e^{2x}\rightarrow du=2e^{2x}dx\rightarrow \frac{1}{2}du=e^{2x}dx\) \(\large \int e^{2x}(1+e^{2x})^3dx=\frac{1}{2}\int u^3du\)